Study Hints Before each class, you should both review your class notes and read the section to be discussed that day. This will allow you to better understand the terminology and concepts discussed in class, to identify concepts that are not clear to you, and to identify questions that you may need to ask during class.

Math 241 is the prerequisite for this course. The review of this material will be brief. Be sure you take all necessary steps to refresh your memory and skills relating to the material in Chapters 1 and 2 of our textbook.

Grading Your grade in this course will be based on your performance on homework, two (2) mid-term exams, and a final exam. The weights assigned to each of these components will be:

Homework  10%
Mid-term exams (3) 60%
Final exam  30%

Course grades will be determined according to the following scale:

A 90 –100
B 80 – 89
C 70 – 79
D 60 – 69
F 0 – 59

Note that the deadline to drop this course with a grade of W is \underline{Thursday, October 2, 1997}.

Exams There will be two (2) exams during the semester. Tentative dates and topics for these exams are:

Wednesday, September 24  Chapters 1, 2, 4, and 5
Wednesday, October 29  Chapters 6 and 7

Make-up exams will be given only for documented reasons of illness, family emergency or participation in a University sponsored event. Excuses such as oversleeping, forgetting the time or location of the exam, and lack of studying are explicitly noted as unacceptable grounds for the administration of a make-up exam.

A comprehensive final will be given at \underline{2:00p.m.} on \underline{Wednesday, December 10, 1997}. (Do not trust exam schedules published by unofficial sources, e.g., the Mortar Board.)

Homework A minimum set of homework problems will be distributed at the beginning of each chapter. You are responsible for working these problems as we work through the book. All problems marked with an asterisk (^{∗}) will be collected and graded once each week (usually on Monday, but this may change).

Your homework average will be determined after dropping your three (3) lowest homework scores \underline{No late homework will be accepted.}

Attendance Regular class attendance is important. Consistent with the USC Undergraduate Bulletin, a grade penalty may be applied to any student missing more than four classes (10%) during the semester.

Academic Honesty Cheating and plagiarism will not be tolerated in this course. You may discuss homework problems with others, but do not copy solutions from another student or from a book. Violations of this policy will be dealt with in a matter consistent with University guidelines.